OpenStudy (anonymous):

MEDAL!!!!!! 56 kd0116 HELP PLEASE 2. Find the gross pay per pay period. Annual salary: $71,290 Pay period: semimonthly (Points : 1) $2970.42 $5940.83 $10,184.29 $11,881.67 Question 3. 3. Brett earns $19.20/h and he gets overtime for hours exceeding 40 per week. How many full hours must he work to earn a gross weekly pay of exactly $1056? (Points : 1) 46 49 50 51

OpenStudy (texaschic101):

gross pay per pay period.....pay period is twice per month...12 months in a year....so that means there is 24 pay periods. annual pay = 71,290......for 24 pay periods = 71,290/24 = $ 2970.42 per pay period ====================== he gets $19.20 per hr -- (nice job).....if he worked 40 hrs, that would be $768 I am assuming he gets paid time and a half for overtime...that means for every hr worked overtime he would make : 19.20/2 + 19.20 = 9.6 + 19.2 = 28.80 per overtime hr. so 1056 - 768 = 288.....288/28.80 = 10 hrs.....so basically he would have to work 50 hrs in order to make 1056. His regular 40 hrs (at 19.20 per hr) + 10 hrs overtime (at 28.80 per hr).

OpenStudy (anonymous):

wow thanks so much!

OpenStudy (texaschic101):

no problem :)